OpenStudy (anonymous):

find all real solutions to 8x^3 + 27 = 0

OpenStudy (anonymous):

try first to factorize your LHS

OpenStudy (anonymous):

(2x-3) (4x-6x+9) I know that part. The thing is seeing if -3/2 is a solution or 3+ or - 3 square root of 8/ 4 is not or a possible solution also

OpenStudy (anonymous):

the factorization is (2x+3)(4x^2-6x+9)=0 which gives 2x+3=0=>x=-3/2 and when you check the discriminant of the second factor you'll find out its negative=> the only solution is x=-3/2

OpenStudy (anonymous):

why does the the other factor (with the quadratic formula) is not a solution?

OpenStudy (anonymous):